SpaceRef

SpaceRef


NOAA GOES SATOPS Morning Report: Wednesday, February 16, 2005

Status Report From: National Oceanic and Atmospheric Administration
Posted: Thursday, February 17, 2005

DAY 47

DMSP

All DMSP operations were nominal over the past 24 hours.

POES

SOCC engineering is continuing to monitor STX-1 on N-17.

N16 rev 22689 at Fairbanks 1224Z: None of the data sets were recoverable from the IMUX when post-pass transfers were attempted. The CDA reported that a significant software problem appears to exist on all 3 antennas, wherein the IMUX does not go into record automatically and possibly cannot be placed into record manually. The CDA is investigating. L2B and L4A will go back into record prior to the next pass and will be lost. G2B was replayed and recovered on the next pass. Attempted to recover L1B on rev 22691 at Fairbanks were unsuccessful. L1B failed to lock on either the Bit Sync or Frame Sync..

All other POES operations were nominal over the past 24 hours.

GOES

GOES spacecraft have started battery reconditioning in preparation for the upcoming eclipse periods. GOES-9 maneuver was successful.

All GOES operations were nominal over the past 24 hours.

Questions or comments to Keith.Amburgey@noaa.gov

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.